Changes between Version 23 and Version 24 of Status


Ignore:
Timestamp:
Nov 13, 2008 1:19:02 PM (11 years ago)
Author:
gaew
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • Status

    v23 v24  
    44
    55Please contact Gary Llewellyn (Science / operations coordinator) if you have any questions or concerns: gaew@nerc.ac.uk (01865 374 391)
     6
     7== ARSF VOCALS Sitrep 11.11.08 ==
     8
     9''Monday’s sortie (VA11) comprised of flying 72 degrees West along the coast to the South of Arica (as far as 23 degrees, 30 minutes South).  This was flown in coordination with the 146. Both aircraft were measured pollution along the coast. Take-off was at 08:24 hrs (local), the flight lasted 4 hours and 234 GB of Eagle and Hawk data were collected.''
     10
     11''Due to the prevailing weather patterns and the expectation of more useful weather and pollution events in following days we will not fly on Tuesday. The day will be spent resting and handling data.''
     12
     13
     14== ARSF VOCALS Sitrep 09.11.08 ==
     15
     16''Sunday’s flight (VA10) was to proceed at approximately 15 000ft from point alpha (20 degrees South, 72 degrees West) to 20 degrees South, 76 degrees West or when contacted by the 146 crew, if sooner. At that point both aircraft were to fly back to point alpha, the Dornier above any cloud and the 146 within. Unfortunately, a problem starting the Dornier’s starboard engine delayed its take-off. The mission scientist on the Dornier contacted the science team at Hotel El Paso and it was decided to continue the flight despite this delay. Take-off was approximately two hours late (11:00 hrs local), the flight lasted 4 hrs and 12 minutes and 300 GB of Eagle and Hawk data were collected.''
     17
     18
     19
     20== ARSF VOCALS Sitrep 08.11.08 ==
     21
     22''A test flight (VA09) was conducted in the afternoon of the 06.11.08. This lasted for 1 hour and 30 minutes during which time ‘just under’ 100 GB of Eagle and Hawk data were collected and the Leosphere LiDAR was demonstrated to be in full working order. ''
     23
     24''Friday (07.11.08) was a change day for the pilots; in the morning David Davies and James Johnson departed Arica, Chile. There was no flying scheduled for the Dornier. The day was used to purchase oxygen and masks, to repair the QC computer (remount drives), to check that backed up data were complete (some files were being lost in transfer when the QC computer was hot or if a drive was near its capacity), evaluation of the storage strategy for those data to be collected during the remaining sorties in the campaign and the continuing process of quality checking data. In the evening Lance Dutton arrived in Arica.''
     25
     26''Saturday was a hard down day. The regular evening meeting with those scientists using the Dornier identified the mission objectives for the following day. Later that evening Carl Joseph returned to Arica. Carl and Lance will fly the Dornier for the remaining campaign in Arica but David Davies will return for the Dornier’s transit back to the UK.''
     27
     28
     29
     30== ARSF VOCALS Sitrep 06.11.08 ==
     31
     32''The sortie yesterday (05.11.10) commenced with a 10:00 hrs (local) take-off. The flight followed the coast 250 nautical miles to the South with the aim of finding and mapping aerosol concentrations. The outward leg was flown in a saw-tooth manner with a maximum altitude of 15 000ft, the return leg was at 8500 ft. The flight lasted for 3 hrs and 11 minutes and 222.6 GB of Eagle and Hawk date were collected.''
     33
     34''Yesterday afternoon the Leosphere LiDAR was removed from the Dornier. Today a technical expert from its makers has arrived to render repairs or at least facilitate an improved functionality. Should this be successful data are intended to be collected this afternoon. One of the drives in the QC computer has failed but a solution has been devised that will allow data to be downloaded, backed-up and quality checked for the rest of the campaign.''
     35
     36
     37
     38== ARSF VOCALS Sitrep 04.11.08 ==
     39
     40''On Monday the Dornier took off, flew at 10:00 hrs (local) and interception with the flight track of the 146 at 11:00 hrs. The main objective of the flight (VA06) was to map and characterise the high level aerosol concentrations (between 11 000 and 13 000ft) in order to determine the pollution gradient within the selected area. The flight lasted for 3 hrs and 26 minutes and 200 GB of Eagle and Hawk date were collected.''
     41
     42''Other than the Leosphere LiDAR, which has improvised modifications to stabilise it, instruments and aircraft are performing well.''
     43
     44''Take-off this morning (Tuesday) was at 08:30 hrs (local). During the sortie (VA07) the Dornier climbed to 15, 000 ft and proceeded to point alpha (20 degrees South, 72 degrees West). From point alpha the Dornier continued to 20 degrees South, 76 degrees West and returned at 10 000ft (in the most relevant aerosol layer). This is the Dornier’s standard flight configuration for the VOCALS campaign (e.g. VA03). The flight lasted for 4 hrs and 12 minutes and 278.5 GB of Eagle and Hawk date were collected.'' 
     45
     46
     47
     48== ARSF VOCALS Sitrep 02.11.08 ==
     49
     50''Take-off on Sunday the 2nd was at 09:00 hrs (local). The flight (VA05/VO05/VOC05) consisted of a ‘saw-tooth profiles’ where on four occasions the Dornier ascended to 15, 000ft and then descended to 4000ft. A fifth decent was in a spiral at 500 ft / minute over homogeneous cloud; the allowed the Berlin instrument to gain a 360 degree view angle over a range of altitudes.  The objective was to monitor aerosols and the flight took 3 hours and twenty minutes during which 270 GB of Eagle and Hawk data were collected. An improvised temporary fix of the Leosphere LiDAR allowed some data to be collected but this instrument still required attention from its manufacturer. The aircraft and other instruments (GRIM, PSASP, AIMMS, Berlin polarimeter, AW spectrometer) continue to perform well.''
     51
     52''We are aware and regularly informed of the worsening situation in Peru and this is information is included in operations planning. The availability of on board oxygen continues to be an issue and may require extra supplies later in the week. However, a very well received BBQ and social get-together was hosted at the El Paso Hotel. This provided a good opportunity to informally discuss the progress of the campaign, research interests and our future flights.''
     53
     54
     55
     56
     57== ARSF VOCALS Sitrep 01.11.08 ==
     58
     59''Our third science flight (VA03) was flown on the Thursday the 30th (October). It comprised of a science flight to the Ron Brown US survey vessel and then four transects in the direction of the Ron Brown’s radar. The total flight time was 4 hour and 20 minute. This was the first of our flights with the under-wing pylons fitted. Unfortunately, the AIMMS 20 probe (fitted in one of the right pods) failed to store data; this problem was resolved soon after landing. The Leosphere LiDAR is still malfunctioning but it is hoped that discussions with the manufacturer will render it serviceable soon. The new system disks for the Specim Eagle and Hawk performed very well and 251 GB were collected.''
     60
     61''As data are collected at 15, 000 ft oxygen usage has been high. Oxygen was therefore replenished in Iququie.''
     62
     63''VA04 was flown yesterday (Friday). As with VA03, take off was at 08:00 hrs (local time). The Dornier was taken to 15, 000 ft and proceeded to point alpha (20 degrees South, 72 degrees West). Point alpha is the standard science start point though data are also collected during assent and decent.  From point alpha the Dornier continued to 20 degrees South, 76 degrees West and returned. The flight time was four hours during which 270 GB of Specim data were collected and included minor saturation in a small section of the Eagle data (~  530nm), thus providing important information regarding instrument settings in this environment.''
     64
     65
     66 
     67== ARSF VOCALS Sitrep 28.10.08 ==
     68
     69''A second ARSF flight was conducted today. The Eagle and Hawk performed well but the Eagle continued to demonstrate disk instability at high altitude. We hope that this will soon be resolved once the hermetically sealed system disks (presently held by customs) are fitted. The flight was a test flight for the ‘Berlin polarometer’ and the Leosphere due to concerns regarding the laser. Contacts are being made to extract local GPS data collected as part of the IGP/DGF network and run by the Centre for Seismology at the University of Tarapaca. This will allow use to better process those data collected on the 26th and today and may mitigate the immediate deficiency in our capability to collect GPS data (this will be resolved on delivery of the GPS receiver tripod). The aircraft itself continues to operate without fault or problems.''
     70
     71
    672
    773== 22nd September ==